# Skeleton: A New Framework for Accelerating Language Models via Task Neuron Localized Prompt Tuning

### v2

@@ 4.1 Quantifying Task Relevance of Neurons
This section describes the process of computing the task relevance of each neuron. We adopt an attribution method (Yang et al., 2023, 2024) to quantify the skill relevance of neurons in the pre-trained language models. Formally, given a task-specific prompt tokens $P_{\tau}$ and text input tokens $x$ provided to $\mathcal{P}_{\theta}$, the contribution of the $i$-th neuron to the prediction of an output $y$ is defined as follows:
$A^{(P_{\tau},y,x)}_{i}(h)=|h_{i}\times\frac{\partial\mathcal{P}_{\theta}(y|[P_{\tau};x])}{\partial h_{i}}|$ (2)
@@
where $h$ corresponds to a hidden representation, and $\partial\mathcal{P}_{\theta}(y|[P_{\tau};x])/\partial h_{i}$ means the gradient of the logit $\mathcal{P}_{\theta}(y|[P_{\tau};x])$ with respect to the $i$-th neuron.
@@
4.2 Identifying Task-relevant Neurons.
@@
4.3 Task Neuron Localization.
@@
4.4 Task Neuron Localized Prompt Tuning
Our Skeleton framework follows the four steps to train a model: (1) the first training step for injecting prior task knowledge; for injecting prior task knowledge; (2) the quantification step of computing task relevance for each neuron; (3) the task neuron localization step; (4) the final prompt tuning step for condensing task knowledge.
The Prior Knowledge Injection Step.
The attribution method quantifies the contribution of each neuron to the prediction of an output; thus, a language model must have the ability to solve a specific task in order to accurately quantify the task-specific knowledge of each neuron. Therefore, we first execute the prior knowledge injection process. Specifically, we train a language model $\mathcal{P}_{\theta}$ via prompt tuning and get a task-specific prompt $P_{\tau}^{\prime}$.
The Quantification Step.
After injecting task knowledge into the language model, we quantify the task relevance $A_{i}$ of each neuron for $\mathcal{P}_{\theta}$ and $P_{\tau}^{\prime}$ using Equation 2 and 3.
The Task Neuron Localization Step.
After quantifying the task relevance $A_{i}$, we first sort the neurons by the task relevance in descending order. Then, we select and maintain only the top-$p$ neurons $\theta_{\tau}\subset\theta$ from the language model by pruning task-irrelevant neurons (Section 4.3). Note that we use a Binary search algorithm to find the optimal ratio of neurons $p\in[0.0,0.05,...,0.95,1.0]$. This search process requires only a maximum of five trials since we search pruning ratio in 0.05 units.
The Knowledge Condensation Step.
During the task neuron localization step, we condense the task knowledge into only the task neurons. Although we never update the original parameters of the language model, we can condense the task knowledge into the task neurons by updating only $P_{\tau}^{\prime}$. We train the task neuron localized model $P_{\theta_{\tau}}$ and get a new knowledge-condensed task prompt $P_{\tau}^{\prime\prime}$. The Binary search algorithm determines the optimal pruning rate after this step.
Algorithm 1 Skeleton Framework
@@ 4.4 Task Neuron Localized Prompt Tuning
2:a task neuron localized model $\mathcal{P}^{*}_{\theta_{\tau}}$
3:train $\mathcal{P}_{\theta}$ via prompt tuning using $P_{\tau}$ and get $P^{\prime}_{\tau}$
@@
6:sort target neurons in descending order by $A_{i}$
@@
13:     get top-$p$ task neuron localized model $\mathcal{P}_{\theta_{\tau}}$
14:     train $\mathcal{P}_{\theta_{\tau}}$ via prompt tuning using $P^{\prime}_{\tau}$ and get $P^{\prime\prime}_{\tau}$
@@
21:return $\mathcal{P}_{\theta_{\tau}}^{*}$
Model Method SST-2 IMDB AGNews MRPC CB
BERT (110M) Fine-tuning (Single) 92.39 92.50 94.12 82.35 78.57
Fine-tuning (Multi) 91.36 92.16 93.98 78.02 79.76
Prompt-tuning 89.36 88.56 90.58 67.07 69.63
Skeleton 89.21 88.65 89.94 68.05 69.04
@@
RoBERTa (125M) Fine-tuning (Single) 94.34 94.56 94.61 85.13 86.90
Fine-tuning (Multi) 94.65 94.70 94.64 86.85 72.02
Prompt-tuning 93.84 92.72 91.87 67.23 66.06
Skeleton 92.81 91.50 90.79 67.89 69.04
@@
Table 1: Task Neuron Localized Prompt Tuning Results. We report the accuracy of five datasets for Skeleton and other baselines. We train models using each method by three trials and report the averaged accuracy. Pruning Rate (%) and Speed up ($\times$) mean the ratio of neurons deactivated after applying our method and the inference speed up compared to the Prompt-tuning, respectively. The values in parenthesis are the number of pruned parameters. The pruning rates are calculated based on the maximum number of prunable parameters.
4.5 Details of the Skeleton Framework
Satisfying Dimension Integrity.
This section describes the detailed task localization processes; the entire process is shown in Figure 2. We prune task-irrelevant neurons from a language model. Recent language models have been trained using the transformer architecture, but they consist of complex neural computation processes. For example, sequentially connected networks are dependent on each other in terms of dimensionality. Suppose there are two sequential feed-forward networks (FFNs), $W_{1}\in\mathbb{R}^{d\times m}$ and $W_{2}\in\mathbb{R}^{m\times d}$. If we prune the first FFN and get a pruned weight $\tilde{W}_{1}\in\mathbb{R}^{d\times m^{\prime}}$, this pruning process affects the inconsistency of the dimension with the weight in the second FFN. Specifically, $W_{2}$ is forced to be pruned as $\tilde{W}_{2}\in\mathbb{R}^{m^{\prime}\times d}$. Therefore, we implement the task neuron localization algorithm by reflecting the dimension dependency of two FFNs. In addition, suppose the representation computed by the second FFN, $v\in\mathbb{R}^{d}$, is connected to the representation of a residual connection network, $u\in\mathbb{R}^{d}$. Then, if we prune the second FFN and get a pruned weight $\tilde{W}_{2}\in\mathbb{R}^{m\times d^{\prime}}$, this process also triggers the inconsistency of the dimension with the representation transferred by the residual connection, making $v\in\mathbb{R}^{d^{\prime}}$. Thus, we insert zero values into the pruned dimension of the second FFN’s representation, $v$, to ensure its integrity.
Accelerating Inference Speed.
Our framework does not just convert the pruned value of weights to zero values but eliminates the entire column (i.e., neuron) of them. Therefore, our framework accelerates the language model inference speed. We do not consider other types of networks (e.g., Attention modules, Embedding layer, Language model head) since they play a crucial role in the model’s language understanding.
Furthermore, the parameters of FFNs constitute most of the prunable parameters (about more than $65\%$); thus, pruning even only the FFNs is sufficient. The experimental results for other modules are described in Section 5.4.
5 Experiments
5.1 Experimental Setup
Datasets.
We conduct experiments on five widely-used natural language understanding datasets [^1] https://huggingface.co/docs/datasets/ (Maas et al., 2011; Zhang et al., 2015; Wang et al., 2018). Specifically, we utilize SST-2, IMDB (sentiment analysis); AGNews (topic classification); MRPC (semantic textual matching); CB (natural language inference) to verify the applicability of our method. We split the train set and use 10% of them as a validation set.
Models and Training Methods.
@@ 5.1 Experimental Setup
We utilize two transformer variants  [^2] https://huggingface.co/docs/transformers/, BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), for our experiments since they are some of the most popular basic language models. We select a prompt tuning method (Lester et al., 2021) to compare with our method, and also include the results of basic fine-tuning and multi-task fine-tuning to show the upper-bound performance of each task.
Implementation Details.
@@
We evaluate Skeleton and baselines on NVIDIA A5000 GPU. We train models using $l=20$ prompt tokens during 100 epochs with early-stop conditions for all datasets. In addition, we set the margin $\psi=1.0\%$ for the validation accuracy score used in the Binary search algorithm. We search $p\in[0.0,0.05,...,0.95,1.0]$ to find the best pruning rate in the Binary search algorithm. We quantify the task relevance using only $n=20$ data instances of each training dataset, following Yang et al. (2024). We specify the target modules for task neuron localization as FFN networks since the experimental results for FFN networks outperform those of other modules, shown in Figure 6. We use learning rates $\gamma_{1}\in$ [1e-5, 5e-5] and $\gamma_{2}\in$ [1e-4, 5e-3] for fine-tuning and prompt-tuning, respectively. We use a masked language model probing network including an instruction for solving a task to utilize the pre-trained knowledge of BERT and RoBERTa, as shown in Appendix A.
5.2 Skeleton is Efficient and Robust
Datasets Experiments.
@@ 5.2 Skeleton is Efficient and Robust
We evaluate and report the task performance and efficiency of each method in Table 1 for the five natural language understanding benchmarks. We adopt BERT-base and RoBERTa-base to conduct the experiments. The results show that our method performs comparably to the prompt tuning method and even outperforms it in accuracy for some datasets using only a small number of task-localized parameters (e.g., at most 65% of pruning rates). Surprisingly, the pruned networks show at most ×$\times 1.72$ speed up; thus, these results reveal that our method successfully localizes efficient task-relevant sub-networks from language models. We describe the detailed elapsed time (sec) of Skeleton models and prompt tuned model in Figure 3. These results reveal that our method significantly outperforms the prompt tuning method in inference speed.
@@
Figure 3: Elapsed Time and Inference Speed-up. We plot the elapsed time and inference speed-up for BERT-base and RoBERTa-base on the four datasets. The values above the bar plot are the speed-up ratios.
Hyperparameter Experiments.
@@
We conduct experiments to investigate the time efficiency of our framework for varying hyperparameters, as shown in Figure 4. Specifically, we search varying batch sizes and the number of input tokens to demonstrate efficiency by comparing our framework to the prompt tuning method. We sample random texts with $k$ tokens and input them into the BERT-base model for the experiments. For the batch size experiments, we evaluate the inference speed with varying batch sizes $\beta\in[16,32,64,256,512]$, fixing the number of tokens at $64$. For the number of tokens experiments, we measure the inference speed with multiple numbers of tokens $k\in[32,64,256,512]$, using the batch size of $64$. The results reveal that our framework shows robust efficiency improvement in various hyperparameters, showing better speed in higher batch sizes and a widely used number of tokens.
@@
Figure 4: Inference Speed Experiments for Varying Hyper-parameters (Batch size & # of Tokens). We report the results for the pruning rates of 30% and 50% for BERT-base to ensure a fair comparison.
5.3 Task Localization Methods Experiments
We demonstrate the task localization performance of those baselines to justify our selection of the task localization method, Attribution, as shown in Figure 5. Specifically, we select the BERT-base model for the SST-2 and IMDB datasets for the experiments. We prune the top 50% of task neurons by using those methods for a fair comparison. Surprisingly, even Activation and Gradient do not show competitive performance compared to Random. However, Skeleton (i.e., Attribution) outperforms other methods in the two datasets, revealing the excellence of our method.
Figure 5: Task Localization Methods Experiments. We plot the mean accuracy (± one standard deviation for three trials) of BERT-base for SST-2 and IMDB.
5.4 Module-specific Task-localization
Transformer variants have various types of modules (i.g., attention modules and FFNs). Therefore, we examine which module is suitable for maintaining the knowledge of a language model (BERT-base). Specifically, we categorize layers into four segments: (1) All layers (All), (2) Attention layers (Attn), (3) Dense layers (Dense), and (4) Feed-forward layers (FFN). All includes all linear layers. Attn consists of Query (Q), Key (K), Value (V), and Output (O) layers of the attention module. Dense contains $O$ of the attention module and two FFNs. FFN includes two FFNs. We prune the top 30% of task neurons from each module for a fair comparison. Figure 6 represents the results of our method for various types of modules. From these experiments, we reveal that FFNs are suitable for maintaining the knowledge of a language model in our prompt tuning setting. Adopting the FFN module yields faster performance than other modules without compromising accuracy.
Figure 6: Module-specific Task-localization Results. We plot the relationship between the accuracy and speed up of each module, where each diagram corresponds to the categorized module. The red-dotted line is the baseline performance (Prompt-tuning).
5.5 Prior Knowledge Injection is Important
Figure 7: Prior Knowledge Injection is Important. We plot the mean accuracy (± one standard deviation for three trials) of BERT-base for SST-2 and IMDB.
This section describes the importance of injecting prior task knowledge (described Section 4.4) to accurately quantify the task relevance, as shown in Figure 7. We prune the top 50% of task neurons by using two methods (i.e., Skeleton and w/o Prior Task Knowledge) for a fair comparison. The experimental results reveal that prior task knowledge injection has a crucial role in quantifying the task relevance. Excluding the Prior Knowledge Injection step results in slower convergence and lower task accuracy compared to our method.
5.6 Skeleton is Scalable to a Larger Model
@@
We conduct experiments to verify that our method is scalable to a larger model, as shown in Table 2. Specifically, we adopt BERT-large to show the scalability of our framework. These results reveal that our method is scalable to a larger model.
@@
BERT-base (110M) Fine-tuning (Single) 92.39 92.50
@@
BERT-large (340M) Fine-tuning (Single) 92.77 93.57
@@
Table 2: Experiments on various sizes of BERT.
6 Conclusion
In this study, we propose Skeleton, which explores the extreme of the prompt tuning method by localizing task-relevant neurons from the original language model. Skeleton investigates the optimal task-localized sub-network from a language model and condenses task knowledge by utilizing only task-relevant parameters during the prompt-tuning procedure. We demonstrate our method for various benchmarks and reveal that Skeleton dramatically enhances the efficiency of language models, showing significant inference speed boost-up.
Limitations
Our framework accelerates the inference speed of language models, identifying the optimal subnetwork for utilizing language models. However, while our approach exhibited comparable performance and occasionally slightly superior results to the existing prompt tuning method, these improvements were not consistent across overall task performance. In addition, our experiments are limited to five widely-used natural language understanding datasets; thus, we should extend to a wide range of datasets to validate the applicability of our framework. Furthermore, we select feed-forward networks to localize task neurons in language models since they are adequate for inference speed acceleration. However, the entire modules in Transformer should be considered as a localization target to achieve the extreme inference speed up; we leave this point as future research.
Acknowledgements

### v3

@@ 4.1. Task Relevance Quantification
This section describes the process of computing the task relevance of each neuron. We adopt an attribution method (Yang et al., 2024) to quantify the skill relevance of neurons in pre-trained language models. Formally, given a task-specific prompt tokens $P_{\tau}$ and text input tokens $x$ provided to $\mathcal{P}_{\theta}$, the contribution of the $i$-th neuron to the prediction of an output $y$ is defined as follows:
(2) $A^{(P_{\tau},y,x)}_{i}(h)=|h_{i}\times\frac{\partial\mathcal{P}_{\theta}(y|[P_{\tau};x])}{\partial h_{i}}|$
@@
where $h$ corresponds to a hidden representation, and $\partial\mathcal{P}_{\theta}(y|[P_{\tau};x])$ $\partial h_{i}$ means the gradient of the logit $\mathcal{P}_{\theta}(y|[P_{\tau};x])$ with respect to the $i$-th neuron of the representation $h$.
@@
4.2. Task Expert Identification
@@
4.3. Task Expert Localization
@@
4.4. The Entire Process of DoE
Our DoE framework follows the three steps: (1) the quantification step of computing task relevance for each neuron; (2) the task expert determination and localization step; (3) the final prompt tuning step for condensing task knowledge. The algorithm for the entire process is described in Algorithm 1.
(1) The Knowledge Quantification Step.  The attribution method quantifies the contribution of each neuron to the prediction of an output; thus, a language model must have the ability to solve a task in order to quantify the task-specific knowledge of each neuron accurately. Therefore, we first execute the prompt tuning to align the knowledge in the language model. Specifically, we conduct the prompt tuning using a frozen language model $\mathcal{P}_{\theta}$ and get an updated task-specific prompt $P_{\tau}^{\prime}$. After injecting task knowledge into the language model, we quantify the task relevance $A_{i}$ of each neuron for $\mathcal{P}_{\theta}$ and $P_{\tau}^{\prime}$ using Equation 2 and 3.
(2) The Task Expert Localization Step.  After quantifying the task relevance $A_{i}$, we first sort the neurons by the task relevance in descending order. Then, we select and maintain only the top-$p$ neurons $\theta_{\tau}\subset\theta$ from the language model by pruning task-irrelevant neurons (Section 4.3). Note that we use a Binary search algorithm to find the optimal ratio of neurons $p\in[0.0,0.05,...,0.95,1.0]$. This search process requires at most five trials, as the pruning ratio is explored in increments of 0.05.
(3) The Knowledge Condensation Step.  During the task expert localization step, we condense the task knowledge to only the task expert. Although we never update the original parameters of the language model, we can condense the task knowledge into the task expert by updating only $P_{\tau}^{\prime}$. We train the task neuron localized model $\mathcal{P}_{\theta_{\tau}}$ and get a new knowledge-condensed task prompt $P_{\tau}^{\prime\prime}$. The Binary search algorithm determines the optimal pruning rate after this step.
Algorithm 1 DoE Framework
@@ 4.4. The Entire Process of DoE
2:a task-localized model $\mathcal{P}^{*}_{\theta_{\tau}}$
3:update $P_{\tau}$ via prompt tuning using $\mathcal{P}_{\theta}$ and get $P^{\prime}_{\tau}$
@@
6:sort neurons in descending order by $A_{i}$
@@
13:  get top-$p$ task-localized model $\mathcal{P}_{\theta_{\tau}}$
14:  update $P^{\prime}_{\tau}$ via prompt tuning using $\mathcal{P}_{\theta_{\tau}}$ and get $P^{\prime\prime}_{\tau}$
@@
21:  end if
22:end while
23:return $\mathcal{P}_{\theta_{\tau}}^{*}$
4.5. Details of the DoE Framework
Satisfying Dimension Integrity.  This section describes the detailed task expert localization processes; the entire process is shown in Figure 2. We prune task-irrelevant neurons from a language model. Recent language models have been trained using the transformer architecture, but they consist of complex neural computation processes. For example, sequentially connected networks are dependent on each other in terms of dimensionality. Suppose there are two sequential feed-forward networks (FFNs), $W_{1}\in\mathbb{R}^{d\times m}$ and $W_{2}\in\mathbb{R}^{m\times d}$. If we prune the first FFN and get a pruned weight $\tilde{W}_{1}\in\mathbb{R}^{d\times m^{\prime}}$, this pruning process affects the inconsistency of the dimension with the weight in the second FFN. Specifically, $W_{2}$ is forced to be pruned as $\tilde{W}_{2}\in\mathbb{R}^{m^{\prime}\times d}$. Therefore, we implement the task expert localization algorithm by reflecting the dimension dependency of two FFNs. In addition, suppose the representation computed by the second FFN, $v\in\mathbb{R}^{d}$, is connected to the representation of a residual connection network, $u\in\mathbb{R}^{d}$. Then, if we prune the second FFN and get a pruned weight $\tilde{W}_{2}\in\mathbb{R}^{m\times d^{\prime}}$, this process also triggers the inconsistency of the dimension with the representation transferred by the residual connection, making $v\in\mathbb{R}^{d^{\prime}}$. Thus, we insert zero values into the pruned dimension of the second FFN’s representation, $v$, to ensure its integrity. If we do not insert zero values, we can not fully utilize the knowledge of the residual connection.
Accelerating Inference Speed.  Our framework does not just convert the pruned value of weights to zero values but eliminates the entire column (i.e., neuron) of them. Therefore, our framework accelerates the language model inference speed. We do not consider other types of networks (e.g., Attention modules, Embedding layer, Language model head) since they play a crucial role in the model’s language understanding. Furthermore, the parameters of FFNs constitute most of the prunable parameters (about $70\%$); thus, pruning even only the FFNs is sufficient. The experimental results for other modules are described in Section 5.4.
5. Experiments
5.1. Experimental Setup
Datasets.  We suppose that language models are given requests about five widely-used natural language understanding datasets [^1] https://huggingface.co/docs/datasets/ (Maas et al., 2011; Zhang et al., 2015; Wang et al., 2018) from users. Specifically, we utilize SST-2, IMDB (sentiment analysis); AGNews (topic classification); MRPC (semantic textual matching); CB (natural language inference) to verify the applicability of our method. We split the train sets and use 10% of them as a validation set.
Model Method SST-2 IMDB AGNews MRPC CB Avg
BERT
Fine-tuning (Single) 92.39 92.50 94.12 82.35 78.57 87.98
Fine-tuning (Multi) 91.36 92.16 93.98 78.02 79.76 87.05
Prompt-tuning 89.36 88.56 90.58 67.07 69.63 81.04
DoE (ours) 89.21 88.65 89.94 68.05 69.04 80.98
@@ 5.1. Experimental Setup
RoBERTa
Fine-tuning (Single) 94.34 94.56 94.61 85.13 86.90 91.10
Fine-tuning (Multi) 94.65 94.70 94.64 86.85 72.02 88.57
Prompt-tuning 93.84 92.72 91.87 67.23 66.06 82.34
DoE (ours) 92.81 91.50 90.79 67.89 69.04 82.40
@@
Table 1. Decomposition of Experts Results. We report the accuracy of five datasets for DoE and other baselines, such as two fine-tuning methods (upper bounds) and prompt-tuning. We train models using each method by three trials and report the averaged accuracy. Our method, DoE, achieves higher speed-ups (up to $\times$1.73) than other methods by leveraging high pruning rates (up to 65%), while maintaining comparable performance to the existing prompt tuning method.
@@
Models and Training Methods.  We utilize two transformer variants [^2] https://huggingface.co/docs/transformers/, BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), for our experiments since they are some of the most popular basic language models. We select a prompt tuning method (Lester et al., 2021) as the backbone and baseline of our method since it is one of the basic PEFT methods. We note that the typical LoRA (Hu et al., 2021) method is not appropriate for DoE, as it adds low-rank updates to the original weights at inference time, making parameter pruning difficult. Additionally, we also include the results of basic fine-tuning and multi-task fine-tuning to show the upper-bound performance of each task.
@@
Implementation Details.  We evaluate DoE and baselines on NVIDIA A5000 GPU. We train models using $l=20$ prompt tokens during 100 epochs with early-stop conditions for all datasets. In addition, we set the margin $\psi=1.0\%p$ for the validation accuracy score used in the Binary search algorithm. We search $p\in[0.0,0.05,...,0.95,1.0]$ to find the best pruning rate in the Binary search algorithm. We quantify the task relevance using only $n=20$ data instances of each training dataset, following Yang et al. (2024). We specify the target modules for task expert localization as FFN networks since the experimental results for FFN networks outperform those of other modules (Described in Section 5.4). We use learning rates $\gamma_{1}\in$ [1e-5, 5e-5] and $\gamma_{2}\in$ [1e-4, 5e-3] for fine-tuning and prompt-tuning, respectively. We predict output texts using a masked language model probing network for the prompts that include an instruction for solving a task, as shown in Table 3. Specifically, the text templates consist of an instruction and an input text. For the prediction, we use the representation of the mask token (i.e., “<<mask>>”).
5.2. DoE is Efficient and Robust
@@
Unplug-and-play Experiments.  We evaluate and report the task performance and efficiency of the baselines in Table 1 for the five natural language understanding benchmarks. We adopt BERT-base and RoBERTa-base to conduct the experiments. The results show that our method performs comparably to the prompt tuning method, which utilizes the full model architecture, and even outperforms it in accuracy for some datasets using only a small number of task-localized parameters (e.g., at most 65% pruning rates). Surprisingly, the pruned networks show at most $\times 1.72$ speed up; thus, these results reveal that our method successfully localizes efficient task-relevant sub-networks from language models. We describe the detailed elapsed time (sec) of DoE models and prompt tuned model in Figure 3. These reveal that our method significantly outperforms the prompt tuning method in inference speed.
@@
Figure 3. Elapsed Time and Inference Speed-up. We plot the elapsed time and inference speed-up for BERT-base and RoBERTa-base on the four datasets. The values above the bar plot are the speed-up ratios.
@@
Hyperparameter Experiments. We conduct experiments to investigate the time efficiency of our framework for varying hyperparameters, as shown in Figure 4. Specifically, we search varying batch sizes and the number of input tokens to demonstrate efficiency by comparing our framework to the prompt tuning method. We sample random texts with $k$ tokens and input them into the BERT-base model for the experiments. For the batch size experiments, we evaluate the inference speed with varying batch sizes $\beta\in[16,32,64,256,512]$, fixing the number of tokens at $64$. For the number of tokens experiments, we measure the inference speed with multiple numbers of tokens $k\in[32,64,256,512]$, using the batch size of $64$. We conduct experiments on the fixed pruning rates $30\%$ and $50\%$ for the comparison. The results reveal that our framework shows robust efficiency improvement in various hyperparameters, showing better speed in higher batch sizes and a widely used number of tokens.
@@
Figure 4. Inference Speed Experiments for Varying Hyper-parameters (Batch size & # of Tokens). We report the results for the pruning rates of 30% and 50% for BERT-base to ensure a fair comparison.
@@
Experiments on various sizes of BERT. We conduct experiments to evaluate the scalability of our method to larger models, as shown in Table 2. Specifically, we adopt BERT-large to show the scalability of our framework. These results reveal the possibility that our method can be scaled to a larger model.
@@
BERT-base Fine-tuning (Single) 92.39 92.50
@@
BERT-large Fine-tuning (Single) 92.77 93.57
@@
Table 2. Experiments on various sizes of BERT.
5.3. Task Localization Methods Comparison
We demonstrate the task localization performance of those baselines to justify our selection of the task localization method, Attribution, as shown in Figure 5. Specifically, we select the BERT-base model for the SST-2 and IMDB datasets for the experiments. We prune the top 50% of task neurons by using those methods for a fair comparison. Surprisingly, even Activation and Gradient do not show competitive performance compared to Random. However, DoE (i.e., Attribution) outperforms other methods in the two datasets, revealing the excellence of our method.
Figure 5. Task Localization Methods Experiments. We plot the mean accuracy (± one standard deviation for three trials) of BERT-base for SST-2 and IMDB.
Dataset Template Label
SST-2 Text: {text}. The sentiment of the text is <<mask>>. positive / negative
IMDB Text: {text}. The sentiment of the text is <<mask>>. positive / negative
AGNews Text: {text}. The topic of the text is <<mask>>. World / Sports / Business / Science
MRPC Text1: {text1}. Text2: {text2}. The two texts are <<mask>>. different / equivalent
CB Premise: {premise}. Hypothesis: {hypothesis}. The premise and hypothesis have a relationship of <<mask>>. implication / contradiction / neutrality
Table 3. Instruction Templates and Labels for Each Dataset. The templates show how text, premise, hypothesis, and mask tokens are used in each dataset.
5.4. Module-specific Task-localization
Transformer variants have various types of modules (e.g., attention modules and FFNs). Therefore, we examine which module is suitable for maintaining the knowledge of a language model (BERT-base). Specifically, we categorize layers into four segments: (1) All layers (All), (2) Attention layers (Attn), (3) Dense layers (Dense), and (4) Feed-forward layers (FFN). All includes all linear layers. Attn consists of Query (Q), Key (K), Value (V), and Output (O) layers of the attention module. Dense contains $O$ of the attention module and two FFNs. FFN includes two FFNs. We prune the top 30% of task neurons from each module for a fair comparison. Figure 6 represents the results of our method for various types of modules. From these experiments, we reveal that FFNs are suitable for maintaining the knowledge of a language model in our prompt tuning setting. Adopting the FFN module yields faster performance than other modules without compromising accuracy.
Figure 6. Module-specific Task-localization Results. We plot the relationship between the accuracy and speed up of each module, where each diagram corresponds to the categorized module. The red-dotted line is the baseline performance (Prompt-tuning).
5.5. Ablation Study on Knowledge Condensation
We conduct the ablation study to justify the necessity of Knowledge Condensation, as described in The Knowledge Condensation Step (Section 4.4). Specifically, we eliminate neurons the same as the results of the DoE in the ablation study for a fair comparison, and do not conduct the knowledge condensation step. As a result, if we do not use Knowledge Condensation, the performance for SST-2 and IMDB significantly degrades. These results reveal the justification for adopting the knowledge condensation step.
SST-2 IMDB
DoE 89.21 88.65
w/o Knowledge Condense 0.0 48.34
[…]
Table 4. Ablation Studies on Knowledge Condensation.
5.6. Prior Knowledge Aligning is Important
To accurately quantify task relevance, this section highlights the role of aligning prior task knowledge using prompt tuning, as described in The Knowledge Quantification Step (Section 4.4) and illustrated in Figure 7. We prune the top 50% of task neurons by using two methods (i.e., DoE and w/o Prior Task Knowledge) for a fair comparison. The experimental results reveal that prior task knowledge alignment has a crucial role in quantifying the task relevance. Excluding the Prior Knowledge Alignment results in slower convergence and lower task accuracy than our method.
Figure 7. Prior Knowledge Injection is Important. We plot the mean accuracy (± one standard deviation for three trials) of BERT-base for SST-2 and IMDB.
6. Conclusion
In this study, we proposed Decomposition of Experts (DoE), a novel framework that optimizes pre-trained language models by identifying and activating task-specific experts. To the best of our knowledge, our method is the first to propose a novel usage paradigm for language models that temporarily deactivates irrelevant experts (i.e., neurons) and restores the full model to await subsequent tasks. DoE leverages an unplug-and-play strategy to enhance inference efficiency while maintaining task performance. Experimental results show that DoE achieves up to a 1.73x speed-up with a 65% parameter pruning rate across diverse tasks. Comparative evaluations and ablation studies validate the effectiveness of DoE and its components. The scalability of DoE makes it a practical and adaptable solution for transformer-based architectures.
Acknowledgements.